OpenStudy (anonymous):

find measurement of E to the nearest degree DE=8 DF= 6 EF= 10. Ef is the hypotenuse

OpenStudy (dumbcow):

use law of cosines 6^2 = 8^2 + 10^2 - 2*8*10*cosE solve for E

OpenStudy (dumbcow):

If this is a right triangle, you could use simple sin/cos ratios sinE = 6/10 cosE = 8/10